A body of mass 25kg moving at 3ms-1 on rough horizontal floor is brought to rest after sliding through a distance of 2.5m on the floor. Calculate the friction of sliding friction(take g=10ms-2)


1
Expert's answer
2021-05-06T18:47:25-0400

Let's first find the deceleration of the body:


"v^2=v_0^2+2ad,""a=\dfrac{v^2-v_0^2}{2d},""a=\dfrac{0-(3\ \dfrac{m}{s})^2}{2\cdot2.5\ m}=-1.8\ \dfrac{m}{s^2}."

The sign minus means that the body decelerates.

Let's apply the Newton's Second Law of Motion:


"-F_{fr}=ma,""-\mu_smg=ma,""\mu_s=-\dfrac{a}{g}=-\dfrac{-1.8\ \dfrac{m}{s^2}}{10\ \dfrac{m}{s^2}}=0.18"